What is y+3=3/4x in standard form? I know 3 has to be subtracted from each side, making the equation y=3/4-3, but I don't know where to go from here!
Answer by fractalier(6550)

From y+3=(3/4)x
we multiply everything by 4, since fractions aren't part of the standard form, and get
4y + 12 = 3x
Rearranging we get
-3x + 4y = -12